OpenStudy (anonymous):

what is the lcd of 13/36 11/30 and 11/60

OpenStudy (anonymous):

Do you know what a prime number is? Do you know how to factor out a number into primes?

OpenStudy (anonymous):

Uhm is it someething that only goes into it one time?

OpenStudy (anonymous):

Yeah, a prime number something that you can't be factored into other whole numbers except itself and one.

OpenStudy (anonymous):

You want to factor the numbers into primes first.

OpenStudy (anonymous):

 Here is an example for 36.

OpenStudy (anonymous):

So, 36 = 2 * 2 * 3 * 3

OpenStudy (anonymous):

Can you do this for 30 and 60?

OpenStudy (anonymous):

so for 30 its 3*2*5 & 60 its 2*2*3*5 right?

OpenStudy (anonymous):

Yeah, that's right

OpenStudy (anonymous):

 You want to find any time that all three of them have a common factor and cross it out like so.

OpenStudy (anonymous):

got so then how would i get the answer cause it shows me 90 1080 30 180

OpenStudy (anonymous):

Oops, I made a small mistake.

OpenStudy (anonymous):

You have to find lcm for 2 numbers at a time.

OpenStudy (anonymous):

So, lcm(36, 30) = 36 * 30 / (2*3) = 180

OpenStudy (anonymous):

lcm(60, 180) = 180

OpenStudy (anonymous):

so hows wouuld 19/24 29/30 13/60 be?